OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x < -12 or x + 3 greater than or equal to 12 A number line with an open circle on -6, shading to the left, and a closed circle on 9, shading to the right. A number line with a closed circle on -6, shading to the left, and an open circle on 9, shading to the right. A number line with an open circle on -6, a closed circle on 9, and shading in between. A number line with a closed circle on -6, an open circle on 9, and shading in between

jimthompson5910 (jim_thompson5910):

what do you get when you solve 2x < -12 for x?

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

how would you solve 2x < -12 for x

OpenStudy (anonymous):

I don't know

jimthompson5910 (jim_thompson5910):

how would you undo the multiplication of 2

OpenStudy (anonymous):

divion

jimthompson5910 (jim_thompson5910):

division by what

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

so divide both sides by 2

jimthompson5910 (jim_thompson5910):

2x < -12 x < ???

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

2x < -12 x < -6

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how would you graph that?

OpenStudy (anonymous):

The first one

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Solve for x: -3(x - 3) = 21 x = -10 x = 10 x = -4 x = 4

jimthompson5910 (jim_thompson5910):

you could take the traditional route and directly solve for x or you could plug in each answer choice and see which one produces a true equation

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Good job dude!

OpenStudy (anonymous):

oh wait you forgot something

OpenStudy (anonymous):

look again

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

Greg combined different weights of fertilizer, soil, and compost to make 29 kilograms (kg) of potting soil for some plants. The weight of fertilizer was x kg. The weight of soil was 2 kg more than three times the weight of fertilizer, and the weight of compost was half the weight of fertilizer. This is represented in the equation below. x + (3x + 2) + one half x = 29 What was the difference in the weights of soil and compost Greg combined? 3 kg 6 kg 14 kg 17 kg

OpenStudy (anonymous):

I've got this one and 7 more to do and then I'm done

OpenStudy (anonymous):

you got the answer?

OpenStudy (anonymous):

Its the 3rd one

OpenStudy (anonymous):

Good stuff!

OpenStudy (anonymous):

Which ordered pair is a solution to the inequality y > x + 5? (-3, 2) (0, -3) (3, 5) (2, 8)

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

good

OpenStudy (anonymous):

am I getting them right

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - 6y if x = -2, y = 3

OpenStudy (anonymous):

-26 26 -10 10

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is it the second one

OpenStudy (anonymous):

no we have -8-18=-26

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

Slope = y1-y2/x1-x2

OpenStudy (anonymous):

so its the last one

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Third one

OpenStudy (anonymous):

Nope. we have a slope of 3 and we know what y=mx+b is right?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

no hold on lets work through this. y=mx+b where m is slope and b where it crosses the y axis. So we have y=3x+b, but we need to know what b is, but we know it has to cross both of those points, so if we take one of those points and plug the x and y values into the equation and solve for b like this 10=3(1)+b --> 10=3+b ---> b=7 so now we have y=3x+b, which can be rearranged to 3x-y=-7

OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 3x - 18 = 2x - 20 10 - 2x = 4x + 34 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the opposite of the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

Put x's on one side and constants on the other. equation 1: 3x-18=2x-20 ---> 3x=2x-2 ---> x=-2 equation 2: 10-2x=4x+34 ---> -24 -2x =4x---> -24=6x ---> x=-4

OpenStudy (anonymous):

-2>-4

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Evaluate: -8 + 4 -12 -4 4 -32

OpenStudy (anonymous):

4-8=-4

OpenStudy (anonymous):

Simplify quantity 3 times x plus 9 all over 3 3x + 3 x + 3 x + 9 x

OpenStudy (anonymous):

x+3

OpenStudy (anonymous):

\[\frac{ 3x+9 }{ 3}\] the 3x can be divided 3 and the 9 answer\[x+3\]

OpenStudy (anonymous):

forget that one

OpenStudy (anonymous):

Are you sure? I can do it.

OpenStudy (anonymous):

Assume a 150 pound person burns 120 calories after 2 hours of sleep and 240 calories after 4 hours of sleep. Part 1 (2 points): Use calories as the y-coordinate and hours as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 8 hours of sleep? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.#

OpenStudy (anonymous):

(2,240), and (1,120) slope = \[\frac{ y1-y2 }{ x1-x2 } = \frac{ 240-120 }{ 4-2 } = \frac{ 120 }{ 2 } = 60\]

OpenStudy (anonymous):

y=60x+0

OpenStudy (anonymous):

8*60=480

OpenStudy (anonymous):

is that part 1 to 3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

how do I write the part 1

OpenStudy (anonymous):

slope=60

OpenStudy (anonymous):

and part just put y=60x+0 and part just put 8*60=480

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Aldora decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a joining fee of $100 and a monthly fee of $36. If x represents the number of months that Aldora is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time. 100 + 36x = C Aldora has allocated a maximum of $352 to spend on her gym membership. Which number line shows the possible number of months that Aldora can be a member of the gym?

OpenStudy (anonymous):

u said u could do it

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

1

OpenStudy (anonymous):

100+36x=352 solving for x: 36x=252 x=7

OpenStudy (anonymous):

2

OpenStudy (anonymous):

3

OpenStudy (anonymous):

4

OpenStudy (anonymous):

#2

OpenStudy (anonymous):

I have a correction to make

OpenStudy (anonymous):

The question about the soil and compost and stuff the answer is 17kg

OpenStudy (anonymous):

Thank you so much for your help, I'll message u how I did when it gets graded

OpenStudy (anonymous):

Did you make the correction?

OpenStudy (anonymous):

Ya

OpenStudy (texaschic101):

well...I can tell you that you did real good on these questions.....good job both of you :)

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

I hope I did okay, I'm really tired so I could have made a stupid mistake even though this stuff is easy for me. Great job @Kylo744

OpenStudy (texaschic101):

question for ya....can you see smart scores on your computer....because I can't see nothing but solid color color circles

OpenStudy (anonymous):

And thanks @texaschic101

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I got a 70% and that not counting the essay part witch is 6 points

OpenStudy (anonymous):

6 points out of how many?

OpenStudy (anonymous):

one

OpenStudy (texaschic101):

so that means the highest you can get is 76 ?

OpenStudy (anonymous):

ya